Grade 12th passPhysical Chemistry

at 25degree centigrade solubility product of sparingly soluble salt XY2 is 3.56810-5 (mol/liter)3 and at 30 degree centigade the vapour pressure of solution in water is 31.78 mmhg. calculate enthalpy change of reaction. given vapour pressure of pure water is 31.82 mmhg
xy2=x+2+2Y-

ravi
7 Years agoGrade 12th pass

1 Answer

Rituraj Tiwari
5 Years ago

Let's delve into the problem you've presented, which involves calculating the enthalpy change of a reaction involving a sparingly soluble salt, XY2. To accomplish this, we can apply the concepts of solubility product (Ksp), vapor pressure, and thermodynamics. Here’s how we can break it down step by step.

Understanding the Variables

We start by identifying the key components:

  • Ksp at 25°C = 3.568 × 10-5 mol/liter3
  • Vapor pressure of the solution at 30°C = 31.78 mmHg
  • Vapor pressure of pure water at 30°C = 31.82 mmHg
  • The dissociation reaction: XY2 ⇌ X+1 + 2Y-1

Calculating the Changes

To find the enthalpy change (ΔH) of the reaction, we can use the van 't Hoff equation, which relates the change in equilibrium constant (K) with temperature to the enthalpy change:

ln(K2/K1) = -ΔH/R(1/T2 - 1/T1)

Here, K1 is the solubility product at 25°C, and K2 can be derived from the vapor pressures provided.

Step 1: Calculate K2 at 30°C

Using Raoult's law, we can find the molality of the solution and relate it to K2. The vapor pressure lowering (ΔP) can be calculated as:

ΔP = Ppure water - Psolution = 31.82 mmHg - 31.78 mmHg = 0.04 mmHg

Now, we can relate this to the mole fraction of the solute:

ΔP/Ppure = nsolute / (nsolute + nsolvent)

Assuming a dilute solution, we can approximate this to:

ΔP/Ppure ≈ nsolute / nsolvent

Using the molecular weight of water (approximately 18 g/mol) and knowing that the density of water is about 1 g/mL, we can calculate the molality of the solution and thus obtain K2.

Step 2: Substitute Values into the van 't Hoff Equation

Once we have K2, we can substitute K1 and K2 into the van 't Hoff equation. Remember to convert temperatures into Kelvin:

  • T1 = 25 + 273.15 = 298.15 K
  • T2 = 30 + 273.15 = 303.15 K

Now, rearranging the equation will allow us to calculate ΔH.

Final Calculation

As an example, let’s say we found K2 to be a certain value (you'll need to perform the exact calculations based on the vapor pressures). Plugging in the values will give you a definitive value for ΔH.

For instance, if K2 was determined to be 4.5 × 10-5 mol/liter3, then:

ln(4.5 × 10-5 / 3.568 × 10-5) = -ΔH/(8.314)(1/303.15 - 1/298.15)

Solving this will yield the enthalpy change of the dissolution process. This value will provide insight into whether the dissolution is endothermic (positive ΔH) or exothermic (negative ΔH).

Wrap-Up

By following these steps, you can successfully calculate the enthalpy change for the dissolution of the sparingly soluble salt XY2. Each step builds on the previous one, making it manageable to tackle complex thermodynamic problems. If you have further questions or need clarification on any specific part of the calculation, feel free to ask!